A company calculates its expected profits as a function of the quantity of the items it can sell. How much are their expected profits if the company’s profit function is P(q)=q³−2000q+500 and their current sales quantity, q, is 60?

1
Expert's answer
2021-12-23T18:11:02-0500

The expected profit function is given by P(q)=q-2000q+500

To get the expected profit when q=60

Substituting the value of q =60 in the equation.


P(60)=(60)-(2000*60) +500

=216000-120000+500,

P(60)=96500.